7) A turntable with moment of inertia IT = 0.042 kg m2 and radius R = 0.24 m is rotating with ω = 0.68 rad/s. A mouse of mass m = 0.053 kg is standing on the edge of the turntable. a) What are the total moment of inertia and the total angular momentum of this system? b) If the mouse walks towards the center of the turntable, does the angular velocity increase or decrease? Explain the answer conceptually. c) What would the angular velocity be once the mouse reaches the center?
Answer #1

7. (a) I- I t (MR²) = (0.042) +(0.0530.242) II = 0.04505 kg m2 L= Iw and w= 0.68 rad is L=0.0306 kgm24 6 I decreases - we inc